Record of events relevant, and quite rare, about the Tracker

Here should be recorded date/time and details of the following events:

  • Power cycles
  • Tracker power configuration
  • Tracker second step calibration
  • Hardware Failure

Year GMT day action note
2019 031:09:55 Thu Jan 31st set Tracker to conf B  
2019 004:10:14 Fri Jan 04th set Tracker ALL on  
2018 280:12:24 Oct 7th set Tracker to conf B  
2018 243:08:09 Aug 31st set Tracker ALL on  
2018 255:08:54 Aug 13th set Tracker to conf B  
2018 222:13:44 Aug 10th set Tracker to conf A  
2018 213:12:04 Aug 1st set Tracker ALL on  
2018 202:14:25 July 20th set Tracker to conf A there was a power cut on bus 3A at 202:07:52
2017 059:16:32 Feb 28th Power cycle lost S side of ladder HwId 612 TkId -711, tried to power cycle twice, with no joy
2014 272-332 Sept 29th-Nov 28th test with TTCS OFF see
2014 128:12:21 May 8th Power Supply ISS Bus 3A tripped off, Tcrate 0, 2, 4, 5, 6 powered on again at 15:00 ( Ladder HwId 605 TkId -609 with the whole Y-side dead.
2014 059:01:17 Feb 28th Power Supply All Tracker crates powered OFF (T1 off at 01:51) because of TTCE problems, power on at 059:06:10 (see attachment "tracker_temp_when_PA_stop.pdf"

2011 163:xx:xx Jun 12th Power cycle Request from POD to reduce power consumption, therefore Tracker was switched off and on again
2011 202:xx:xx Jul 24th 2nd step cal. Second step calibration with Δt = 220 μs and 235 μs
2011 216:xx:xx Aug 4th 2nd step cal. Due to the new JAP loading permanent status bit were lost, so second step calibration has been performed with Δt = 220 μs and 235 μs
2011 230:10:00 Aug 18th Power supply conf. The power supply on all the crates are now in nominal configuration: all redundant part are on on the HOT side but S9053 that is on on the cold side
2011 236:05:18 Aug 24th 2nd step cal. Second step calibration with Δt = 220 μs and 235 μs
2012 245:21:00 Sep 1st Power supply Due to a failure on ISS power bus 3A, Tcrate 0, 2, 4, 5, 6 were powered off, powered on again at 23:45
2012 285:13:05 Oct 11th Noisy strips HwID 402 TkID +703 had ~ 700 strips tagged noisy (bit 3) instead of usual ~ 30. A re-boot of DSP (even if never detected as DSP error) fixed the issue. This was due to the corruption of parameter 0xD (0xD and 0xE are used during calibration to determine noisy and dead strips, see and was checked looking inside CFG blocks (DT 0x7). Is possible to check these parameters also retrieving parameters with command 2e09.
2012 293:12:20 Oct 19th Power cycle Ladders HwId 302 TkId -814 and HwId 303 TkId -812 with bad calibrations on K side: on TkId -814 noise level of strips from 961 to 1024 (complete 16th VA) increasing rapidly with time, noise level lower than expected on all of remaining strips. An increase of leakage currents and temperature of 1N-4 sensor, near the Hybrid on plane 1, has been also observed. Problem fixed by cycling the power on FE K-side. Due to this (probably a power glitch) VA 12th of HwId 302 TkId -814 and VA 13th of HwId 303 TkId -812 are damaged (well know problem of broken capacitors)
2012 303:09:35 Oct 29th Strip Tag VA 12th of HwId 302 TkId -814 and VA 13th of HwId 303 TkId -812 since damaged (well know problem of broken capacitors) are tagged permanently as bad
2012 306:08:50 Nov 1st Power supply Due to PDS overheating all Tracker crates but T5 are powered off
2013 311:04:39 Nov 7st Power supply Tracker crates T0, T1,T2 and T3 are powered off for AMS power down for 37S docking. FN 060103:see
2011 318:12:10 Nov 14th Tracker power Power off TPD1, TPD3, TPD7 because of PDS overheating, on again at 14:22 and 15:55
2011 319:14:00 Nov 15th Power supply conf. Flight configuration applied on T1, T3 and T7
2011 335:16:00 Dec 1st Power supply Problem with S9051 in crate T3 (K-side of #12,13,16,17,20,21) broken (first cal 1322757335)
2013 345:08:50 Dec 11th Power supply Tracker crates T4, T5,T6 and T7 are powered off to switch all of them from hot to cold digital power, see attached procedure, see LEAD log and TEE log
2012   Feb 15th Strip Tag All bad K-side's (8 ladders -> 3200 strips) tagged as permanently bad. One VA (64 strips) tagged on S-side (first cal 1329271885)
2012   Feb 08th Strip Tag K-side of T3-12,13,16,17,20,21 tagged as permanently bad (first cal 1328720709)
-- GiovanniAmbrosi - 22-Aug-2011
Edit | Attach | Watch | Print version | History: r59 < r58 < r57 < r56 < r55 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r59 - 2020-04-09 - MauraGraziani
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Main All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2020 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding TWiki? Send feedback